OpenStudy (anonymous):

Juliana has created the function f(x) = the quantity of 3x plus 2, divided by 4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that using the inverse function to solve for x = 24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences

OpenStudy (anonymous):

@myininaya @hba

OpenStudy (anonymous):

If you can just explain how to solve a function I can write the sentences myself I'm just terrible at functions

myininaya (myininaya):

Ok, so explain in symbols what f(x) is?

OpenStudy (anonymous):

f(x) is the number of her text, right?

myininaya (myininaya):

No, it says this: f(x) = the quantity of 3x plus 2, divided by 4 to represent the cost of texting on her current plan, where x represents the number of texts It says that f(x) is cost of the texting whereas x is number of texts. I'm just asking you to explain using mathematical symbols what f(x) is?

OpenStudy (anonymous):

isn't it always f(x) = so it's what everything equals?

myininaya (myininaya):

So can you please give it to me in symbols please? I'm just asking you to translate the quantity of 3x plus 2, divided by 4. This is what f(x) is according to your paragraph.

OpenStudy (anonymous):

f(x)= \[\frac{ 3x +2 }{ 4 }\] I don't really understand what you're asking i'm sorry...

myininaya (myininaya):

I was asking for what f(x) was which is the thing you just stated. The equal sign means is.

myininaya (myininaya):

Now they are asking for the inverse function. This means solve for x basically.

myininaya (myininaya):

you have y=(3x+2)/4

OpenStudy (anonymous):

Okay..so what's first

myininaya (myininaya):

Well how do you undo division by 4?

OpenStudy (anonymous):

I really don't know I'm sorry..wouldn't you do something with distributive property I think it's called?

myininaya (myininaya):

\[y=\frac{3x+2}{4}\] \[y=\frac{3x+2}{4 \text{ <--- Do you see this 4 underneath here? } }\]

OpenStudy (anonymous):

Yes, I see the 4

myininaya (myininaya):

Like we need to get that 4 out of the denominator? Since that 4 is in the denominator we recognize this as division by 4. To undo division by 4, we multiply by 4. Whatever we do to one side we do to the other. Because 1/4 * 4=1 Correct?

OpenStudy (anonymous):

Yes, correct.

myininaya (myininaya):

Ok then what do you get after multiplying both sides by 4?

OpenStudy (anonymous):

Do you need to do distributive property? Where you would do 3x * 4 and 2*4 ?

OpenStudy (anonymous):

If so then it would be (12x + 8)

myininaya (myininaya):

Well I'm asking you to multiply 4 on sides....like this 4y=4(3x+2)/4 The whole reason we multiplied by 4 was to undo that division by 4. So now we have 4y=(3x+2) or just simply 4y=3x+2 (parenthesis not needed) Remember your objective is to get x. You are almost there.

OpenStudy (anonymous):

@myininnaya I know this was a long time ago but I am also stuck on this problem and I need some help, would you help me finish the problem???

OpenStudy (anonymous):

f(x)=3x+2/4